# In Search of Cosmic Time: Complete Observables and the Clock Hypothesis

### v1

@@ 5. Finding Cosmic Time
where $H(z)$ is obtained from the cosmological parameters $\Omega_{I},H_{0}$ which are determined by the experimental power spectrum. To be precise, $\tau(z)$ above does not correspond to the proper time of a real observer comoving with the Hubble Flow, since in observational practice, we know the Universe is not perfectly homogeneous and isotropic: not even on cosmologically large scales. The FLRW model is an idealisation to describe the average behaviour of the Universe. The Hubble parameter $H(z)$ in the formula depends on parameters such as $\Omega_{\text{CDM}}$ (dark matter density), which are in part determined using perturbative methods, which take into account small inhomogeneities. Therefore, in measuring the parameter $H(z)$, contributions of inhomogeneity (such as galaxy clusters, voids) and anisotropies are taken into account. Nonetheless, the value of $H(z)$ used in the theoretical calculation of $\tau(z)$ refers to the ‘average’ behaviour of the Universe, described by a homogeneous and isotropic model. We mean that the formula $\tau(z)$ calculating cosmic time is based on a model that idealises the Universe as perfectly homogeneous and isotropic, using an ‘average value’ of $H(z)$. This means that the cosmic time $\tau(z)$ is a global average, which does not reflect local fluctuations or deviations from isotropy, but rather provides an approximated estimate. One could say that, experimentally, the contributions from inhomogeneities and anisotropies of the actual Universe are ‘averaged out to zero’, in the sense that these contributions are present in the data but, when calculating the Hubble parameter $H(z)$, local fluctuations are essentially ‘smoothed out’ and do not significantly influence the overall result. Consequently, experimental measurements approximate the theoretical ideal cosmic time, which refers to an ideally homogeneous and isotropic Universe, analogously to what happens with primary and secondary clocks for measuring UTC. It is the case that $\tau(z)$ that is used to make predictions in cosmology, and it is a partial observable, pace Thiemann (2006).
We thus have that in a more liberalised sense of ‘measurement procedure’, it is the case that we can treat proper time along the Hubble flow as a partial observable in the context of cosmology. This has not involved simply stipulating proper time as partial observable via the clock hypothesis. However, it also has not involved the abandonment of the clock hypothesis altogether. Rather, the clock hypothesis reemerges within cosmology as something like a ‘coordinative definition’ in sense of Reichenbach (1928). That is, the clock hypothesis allows for the coordination of a concept (an ideal clock measuring the Hubble flow temporal parametrisation) with an empirical phenomenon (cosmic time). This broadly logical empiricist understanding accords with other discussions of the clock hypothesis in the recent philosophical literature (Adlam et al. 2022) and makes sense of the physically non-trivial but partially definitional role. [^17] See also Fletcher (2025) who classifies the clock hypothesis as a representational principle. We recover the Friedmann equations and cosmic time whilst keeping both the clock hypothesis and the partial/complete observable distinction, albeit each in modified form. The conservative option for finding cosmic time is thus a live possibility.
6. Summary
@@
In this paper we first provided a brief overview of the complete and partial observables programme as a response to the problem of time that seeks to preserve the Dirac criterion for observables. We then applied recent work on reference frames to disambiguate two important details in the definition of a partial observable and better understand what it means for a physical variable to play the role of a clock in the context of a complete observable. We then considered the status of time in the Friedmann equations. In particular, we reviewed Thiemann’s challenge to the standard interpretation of these equations as dynamical equations. This led to a discussion of his solution in terms of a de-parameterization via a phantom field through the well-known Brown and Kuchař (1995) mechanism. We next considered the idea of the Hubble flow and sought to re-frame comic time as a proper time parameter along the Hubble flow. In this context, we considered the question of the observability and measurability of Hubble parameter and the question of whether we can consider to be a complete observable whose dynamics is described by the unmodified Friedman equations. We then articulated the specific problem of there being no physical system which can be understood to even approximately measure proper time along the Hubble flow. This leads to the final step in the articulation of our proposed solution: the introduction of a more liberalised sense of ‘measurement procedure’ in the context of cosmic time. This more liberalised notion draws upon the model-based account of time measurement developed in the context of atomic clocks and the measurement of Coordinated Universal Time due to Tal (2016). Provided one is willing to reinterpret our criteria of observability in light of the clock hypothesis and the model-based account of measurement one can and preserve the Friedmann equations and find time in cosmology.
@@ 7. Acknowledgements
We are extremely grateful to David Sloan for detailed comments on a draft manuscript. KT would also like to thank Sean Gryb for helpful discussion.
@@ Appendix A Gauge Freedom and Gauge Fixings
The framework of Dirac observables has as its intended goal the removal of underdetermination in phase space dynamics due to the presence of gauge freedom. This is clearly the important point and not whether or not the letter of the Dirac criterion or the spirit of general covariance have been respected. The observables of the theory are required to be gauge-invariant quantities. Here we briefly summarise some formal aspects relating gauge invariance and gauge fixings in order to clarify the formal status of cosmic proper time within the foundations of general relativity. See Gomes (2024) for further details.
@@
Equivalently we can make use of the so-called projection operator $f_{\sigma}:\varphi\rightarrow f_{\sigma}^{*}\varphi$, an equivalent of the section map, which takes any element of a given fibre to the unique image of the section. It is an embedding map, acting within a fibre and it is characterised by the auxiliary condition $F_{\sigma}(\varphi)=0$, making the choice of a reference frame (or a section) analogous to a gauge-fixing procedure. The use of the projection operator $f_{\sigma}:\Phi\rightarrow\Phi$ instead of the section map $\sigma:[\Phi]\rightarrow\Phi$ codifies the external sophistication demand of the unnecessity to represent elements $[\varphi]$ of $[\Phi]$ intrinsically by some parameterization of $[\Phi]$. Thus the discard of the structure-first, or internal approach (Dewar 2019; Jacobs 2021). The quantity resulting from the choice of a section is the relational, gauge-invariant observable $f_{\sigma}^{*}\varphi\equiv(\varphi)_{F}$. See Figure 2.
@@
In order for the cosmic proper time $\tau$ to be considered a physical clock dynamically coupled with the Hubble parameter (and thus with the metric), the following condition must be met: If $(g_{ab},\tau)$ is a dynamically possible model of the theory, then neither $([d^{*}g]_{ab},\tau)$ nor $(g_{ab},d^{*}\tau)$ is, $\forall d\in Diff(\mathcal{M})$. Thus, the choice of $\tau$ as the RRF clock (rather than one of its diffeomorphic copies) provide a unique representation of $H(\tau):=H\circ\tau^{-1}$ for some initial data, which is thus a bona-fide gauge-invariant, complete observable. For this reason, $\tau$ fixes the gauge for the FLRW metric and by definition is such that there is no longer gauge freedom in the theory, and the potential for underdetermination. However, given the correspondence between the choice of a reference frame and the choice of a gauge (Bamonti 2023; Gomes 2024), we always have the possibility of changing reference frames and obtaining new relational observables.

### v2

@@ 5. Finding Cosmic Time
where $H(z)$ is obtained from the cosmological parameters $\Omega_{I},H_{0}$ which are determined by the experimental power spectrum. To be precise, $\tau(z)$ above does not correspond to the proper time of a real observer comoving with the Hubble Flow, since in observational practice, we know the Universe is not perfectly homogeneous and isotropic: not even on cosmologically large scales. The FLRW model is an idealisation to describe the average behaviour of the Universe. The Hubble parameter $H(z)$ in the formula depends on parameters such as $\Omega_{\text{CDM}}$ (dark matter density), which are in part determined using perturbative methods, which take into account small inhomogeneities. Therefore, in measuring the parameter $H(z)$, contributions of inhomogeneity (such as galaxy clusters, voids) and anisotropies are taken into account. Nonetheless, the value of $H(z)$ used in the theoretical calculation of $\tau(z)$ refers to the ‘average’ behaviour of the Universe, described by a homogeneous and isotropic model. We mean that the formula $\tau(z)$ calculating cosmic time is based on a model that idealises the Universe as perfectly homogeneous and isotropic, using an ‘average value’ of $H(z)$. This means that the cosmic time $\tau(z)$ is a global average, which does not reflect local fluctuations or deviations from isotropy, but rather provides an approximated estimate. One could say that, experimentally, the contributions from inhomogeneities and anisotropies of the actual Universe are ‘averaged out to zero’, in the sense that these contributions are present in the data but, when calculating the Hubble parameter $H(z)$, local fluctuations are essentially ‘smoothed out’ and do not significantly influence the overall result. Consequently, experimental measurements approximate the theoretical ideal cosmic time, which refers to an ideally homogeneous and isotropic Universe, analogously to what happens with primary and secondary clocks for measuring UTC. It is the case that $\tau(z)$ that is used to make predictions in cosmology, and it is a partial observable, pace Thiemann (2006). We thus have that in a more liberalised sense of ‘measurement procedure’, it is the case that we can treat proper time along the Hubble flow as a partial observable in the context of cosmology. This has not involved simply stipulating proper time as partial observable via the clock hypothesis. However, it also has not involved the abandonment of the clock hypothesis altogether. Rather, the clock hypothesis reemerges within cosmology as something like a ‘coordinative definition’ in sense of Reichenbach (1928). That is, the clock hypothesis allows for the coordination of a concept (an ideal clock measuring the Hubble flow temporal parametrisation) with an empirical phenomenon (cosmic time). This broadly logical empiricist understanding accords with other discussions of the clock hypothesis in the recent philosophical literature (Adlam et al. 2022) and makes sense of the physically non-trivial but partially definitional role, cf. (Fletcher 2025). We recover the Friedmann equations and cosmic time whilst keeping both the clock hypothesis and the partial/complete observable distinction, albeit each in modified form. The conservative option for finding cosmic time is thus a live possibility.
6. Summary and Outlook
@@
The complete and partial observables programme is a response to the problem of time that seeks to preserve the Dirac criterion for observables. Recent work on reference frames allows us to disambiguate the definition of a partial observable and better understand what it means for a physical variable to play the role of a clock in the context of a complete observable. Thiemann’s approach to the interpretation of the Friedmann equations as dynamical equations leads to their de-parametrisation via a phantom field through the Brown and Kuchař (1995) mechanism. An alternative approach is based on the idea of comic time as a proper time parameter along the Hubble flow. In this context, we can re-consider the observability and measurability of the Hubble parameter and its status as a complete observable whose dynamics is described by the unmodified Friedman equations. The problem is then that there is no physical system which can be understood to even approximately measure proper time along the Hubble flow. This leads to the introduction of a more liberalised sense of ‘measurement procedure’ in the context of cosmic time drawing upon the model-based account of measurement. Provided one is willing to reinterpret our criteria of observability in light of the clock hypothesis and the model-based account of measurement, one can preserve the Friedmann equations and find time in cosmology.
The problem of time is often presented as arcane question in the foundations of quantum gravity. If the loss of time within our physical theories were only an issue at the Planck scale then it would hardly seem an urgent one. We hope to have given good reasons to dispel such complacency. In particular, we have show how aspects of the problem occur even in the familiar context of classical cosmology and even at scales upon which we already have a wealth of observable evidence, such as those relevant to the age of the universe. The way in which we approach the problem can have empirical physical consequences. Moreover, on our account, the cosmic problem of time is closely connected to both important methodological questions relating to idealisation and the nature of measurement and to familiar foundational topics such as the interpretation of the clock hypothesis and the nature of gauge degrees of freedom. As such, we suggest the search for cosmic time warrants more attention from both philosophers of science and scientists alike.
@@ 7. Acknowledgements
We are extremely grateful to David Sloan, two anonymous referees, and an editor for helpful comments on draft manuscript(s). KT would also like to thank Sean Gryb for helpful discussion.
@@ Appendix A Gauge Freedom and Gauge Fixings
The framework of Dirac observables has as its intended goal the removal of underdetermination in phase space dynamics due to the presence of gauge freedom. This is clearly the important point and not whether or not the letter of the Dirac criterion or the spirit of general covariance have been respected. The observables of the theory are required to be gauge-invariant quantities. Here we briefly summarise some formal aspects relating gauge invariance and gauge fixings in order to clarify the formal status of cosmic proper time within the foundations of general relativity. See Gomes2024 for further details.
@@
Equivalently we can make use of the so-called projection operator $f_{\sigma}:\varphi\rightarrow f_{\sigma}^{*}\varphi$, an equivalent of the section map, which takes any element of a given fibre to the unique image of the section. It is an embedding map, acting within a fibre and it is characterised by the auxiliary condition $F_{\sigma}(\varphi)=0$, making the choice of a reference frame (or a section) analogous to a gauge-fixing procedure. The use of the projection operator $f_{\sigma}:\Phi\rightarrow\Phi$ instead of the section map $\sigma:[\Phi]\rightarrow\Phi$ codifies the symmetry-first or external sophistication principle that it is unnecessary to intrinsically represente elements $[\varphi]$ of $[\Phi]$ through a parametrisation of $[\Phi]$. Thus it rejects the structure-first, or internal approach (Dewar2019-DEWSAS-4; jacobs2021symmetries). The quantity resulting from the choice of a section is the relational, gauge-invariant observable $f_{\sigma}^{*}\varphi\equiv(\varphi)_{F}$. See Figure 1.
@@
In order for the cosmic proper time $\tau$ to be considered a physical clock dynamically coupled with the Hubble parameter (and thus with the metric), the following condition must be met: If $(g_{ab},\tau)$ is a dynamically possible model of the theory, then neither $([d^{*}g]_{ab},\tau)$ nor $(g_{ab},d^{*}\tau)$ is, $\forall d\in Diff(\mathcal{M})$. Thus, the choice of $\tau$ as the RRF clock (rather than one of its diffeomorphic copies) provide a unique representation of $H(\tau):=\big{[}\tau^{-1}\big{]}^{*}H$ for some initial data, which is thus a bona-fide gauge-invariant, complete observable. [^26] With the symbol $\big{[}\bullet\big{]}^{*}$, we denote the pullback. For this reason, $\tau$ fixes the gauge for the FLRW metric and by definition is such that there is no longer gauge freedom in the theory, and the potential for underdetermination. However, given the correspondence between the choice of a reference frame and the choice of a gauge (Bamonti, 2023; Gomes2024), we always have the possibility of changing reference frames and obtaining new relational observables.